OpenStudy (anonymous):

PLEASE help me A diameter of a circle has endpoints p(-10,-2) an Q(4,6) a. find the center of the angle. b. find the radius. c. write an equation for the circle.

OpenStudy (er.mohd.amir):

EQ of circle (x-x1)(x-x2)+(y-y1(y-y2)=0 where (x1,y1)=(-10,-2)

OpenStudy (anonymous):

for A i meant to say "find the center of the circle" sorry

OpenStudy (er.mohd.amir):

put values and x2+y2+2gx+2fy+c=0 STD EQ of circle center (-g,-f) and radius=sqrt(g2+f2-c)